The distance between the points (1,3) and (-2,7) is:
25 units
units
7 units
5 units
Each of the remaining questions refer to the parabola y = 2x 2 + 12x + 17
The vertex of the parabola is:
(3,1)
(3,-1)
(-1,3)
(-3, -1)
The focus of the parabola is:
(3, -9/8)
(-3, -9/8)
(-3, -7/8)
(3, -7/8)
The directrix of the parabola is:
The vertical line x = -9/8
The horizontal line y = -9/8
The horizontal line y = 9/8
The horizontal line y = -7/8
The length of the latus rectum of the parabola is:
17 units
1/8 unit
2 units
1/2 unit

Answer by stanbon(75887)

The distance between the points (1,3) and (-2,7) is:
d=sqrt((7-3)^2+(-2-1)^2)=sqrt(16+9)=5
----------------------------
Each of the remaining questions refer to the parabola y = 2x^2 + 12x + 17
The vertex of the parabola is:
Vertex at x=-b/2a=-12/4=-3;y=2(3)^2+12(-3)+17=-1
Vertex at (-3,-1)
-------------------
The focus of the parabola is:
y-17+18 = 2(x^2+6x+9)
(x+3)^2=(1/2)(y+1)
4p=1/2
p=1/8
Focus at (-3,-1+(1/8))=(-3,-7/8)
-----------------
Thehe directrix of the parabola is:
The horizontal line y = -9/8
----------------
The length of the latus rectum of the parabola is:
I think it is 4p=1/2
-----------
Cheers,
Stan H.